OpenStudy (anonymous):

Minh’s data shows that on average only 4.6% of customers were purchasing complementary services available within Media Systems. By using his company’s services, Minh was projecting that these percentages would triple across all user groups within 1 year. a) How many customers would that equate to in total for each group? b) What would be the difference compared to current levels?

OpenStudy (kirbykirby):

Are there numbers associated with this problem? I believe you would have to multiply 4.6 by 3 to get the tripled percentage... but then I don't see numbers to find the number of customers..

OpenStudy (anonymous):

1. The company has 70,000 publications subscribers, 58,000 advertisers, 30,000 telephone services customers, and 18,000 Internet service provider customers.

OpenStudy (kirbykirby):

So, 3*(4.6) will give you the percentage that has tripled. Now, if you covert that percentage as a decimal, and then multiply it to the number for each group of customers, you will get number of people who purchased the complimentary systems

OpenStudy (kirbykirby):

for b) The current levels is the number of customers at the level of 4.6%, so find the number of people at 4.6%. The new levels are the number of customers at the new tripled-percentage. So to find different between current levels: find (# of customers at tripled-percentage) - (# of customers at the level 4.6%)

OpenStudy (anonymous):

I am puzzled because I found the following online and was wondering where the 9.2% comes from. 70000 * 4.60% * 3 = $9660 70000*9.2% =$6440 58000 * 4.60% * 3 = $8004 58000*9.2%=$5336 30000 * 4.60% * 3 = $4140 30000*9.2%=$2760 18000 * 4.60% * 3 = $2484 18000*9.2%=$1656

OpenStudy (anonymous):

OK, if I understand correctly, everything is multiplied by 3 than subtracted by the original number or percent which give us 2, for instance: 4.60 x 3 = 13.8 - 4.60 = 9.2 right? 70000 x .046 = 3220 x 3 = 9660 9660 - 3220 = 6440 right? So in a sense you triple the factors and then subtract 1 and the results that are show are really for two?

OpenStudy (kirbykirby):

Yes they have done question 1) on the left column, and question 2 on the right column. It turns out that finding 70000*(4.6%*3) - 70000*(4.6%) is the same as 70000*(9.2%) since 13.8-4.6 = 9.2

OpenStudy (anonymous):

Okay, thanks, appreciate it.

OpenStudy (kirbykirby):

:) yw